### v2

# FinResearchBench: A Logic Tree based Agent-as-a-Judge Evaluation Framework for Financial Research Agents

Abstract.
Recently, AI agents are rapidly evolving in intelligence and widely used in professional research applications, such as STEM, software development, finance, etc. Among these AI agents, deep research agent is a key category as it can perform long-horizon tasks and solve problems of greater complexity. However, there are few evaluation frameworks and benchmarks that systematically and automatically investigate the capabilities of these research agents. Furthermore, financial research problems have distinct complexity and subtlety. To fill in the gap, we propose FinResearchBench, which is a logic tree based Agent-as-a-Judge and targets specifically for the financial research agents. It provides a comprehensive and automatic assessment of the research agents across 7 key types of tasks in the financial research domain. The contributions of this work are two-folded: (1) the first and innovative Agent-as-a-Judge system that extracts the logic tree of the research outcome and uses it as the intermediate information to present a comprehensive, reliable and robust evaluation; (2) finance oriented that it covers 70 typical financial research questions, spreading across 7 frequently-encountered types of tasks in the domain.
AI agents, financial research agents, logic tree, Agent-as-a-Judge
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generationccs: Applied computing Economics
1. Introduction
Large Language Models (LLMs) have been rapidly developing in recent years and demonstrate superior performance in diverse tasks, such as story creation (Zhao et al., 2023), machine translation (Xu et al., 2024), dialogue system (Yi et al., 2024), sentiment analysis (Zhang et al., 2023b), etc. Furthermore, LLMs are dramatically evolving in both long text understanding and generation (Team et al., 2024; Anthropic, 2024). The intrinsic capability of LLM to generate long text is often limited, which is revealed to be bounded by the training data the model has seen during the pre-training and alignment stages (Bai et al., 2024). Yet, long text generation is essential, especially in professional research areas, such as literature, law, STEM and finance. Numerous methods have been proposed to improve LLM’s competence in long text generation, including scaling the generation length by adding long output training data (Bai et al., 2024), performing generation in a multi-agent framework (Hong et al., 2023), setting up as deep research agents (Schmidgall et al., 2025; Zheng et al., 2025), and so on. Deep research agent is a prominent category of AI agents and is getting more and more important, since it could generate well-articulated, source-attributed research report by autonomously browsing the internet, calling various tools, collecting information from different sources, etc. The research result is of high value and brings huge productivity improvement. Therefore, comprehensive evaluation of these deep research agents becomes critical. Various benchmarks and evaluation methods are proposed, such as HelloBench (Que et al., 2024), Deep Research Bench (Bosse et al., 2025), LongEval (Wu et al., 2025), etc. However, these methods do not fully utilize the results of the intermediate steps, which could be the deciding factor for the quality of the research outcome. Furthermore, existing benchmarks and evaluation methods are of general purpose and face severe challenges when applied for financial research, since it requires deep understanding of finance-specific terminologies, clauses and domain-specific insights. In this work, we propose FinResearchBench, a logic tree based Agent-as-a-Judge that is designed specifically for financial research agents. It first extracts the logic tree from the research outcome and then uses the extracted logic tree to build an agentic system to evaluate the financial research agents. Moreover, FinResearchBench is finance concentrated and covers 7 key types of tasks in the financial research domain. With the aim of the extracted logic tree as the intermediate step, FinResearchBench is able to provide a comprehensive, reliable and robust evaluation for various financial research agents.
Figure 1. The Illustration of Writing Process
2. Related Work
2.1. LLM and Deep Research Agent
Large Language Models (LLMs) have shown remarkable performance in various tasks, such as story creation, machine translation, dialogue system, etc (Zhao et al., 2023; Xu et al., 2024; Yi et al., 2024; Zhang et al., 2023b). In terms of dealing with long text, they also demonstrate superior capability in understanding and processing (Team et al., 2024; Anthropic, 2024). However, the capability of generating long text progresses rather slowly. Many AI agent systems are then developed including single-agent and multi-agent ones. And it represents a paradigm shift from standalone LLMs towards more autonomous, task-oriented systems (Hong et al., 2023; Wu et al., 2023). AI agents largely increase the upper bound of the generation results since they are able to perform more actions, like web browsing, tool calls, code execution, etc. Additionally, AI agents are capable of processing the obtained information more effectively and accurately, such as Chain-of-Throughts (CoT) method (Wei et al., 2022), Tree-of-Thoughts (ToT) method (Yao et al., 2023a), ReAct method (Yao et al., 2023b), etc. Deep research agent is a growing category of AI agents, since it can orchestrate multiple steps in an autonomous manner, including internet browsing, information retrieval, tool calling, etc. It is able to provide in-depth research of highly complicated problems, and thus brings huge productivity improvement and high commercial values.
2.2. General and Domain-Specific Evaluation of Deep Research Agent
The evaluation of AI agents is critical since it determines if the emergent agents are applicable for real-world problems or not. Moreover, the evaluation will guide the direction for further improvement. Numerous general benchmarks and evaluation methods are proposed, such as AgentBench (Liu et al., 2023), ToolQA (Zhuang et al., 2023), HelloBench (Que et al., 2024), DeepResearch Bench (Du et al., 2025), LongEval (Wu et al., 2025), etc. These benchmarks are constructed with diverse tools in different environments, enabling extensive evaluation of given agents. For deep research agent, it is a subcategory of the AI agents and performs multi-step orchestration in an autonomous manner. It produces research report that contains rich information and is well articulated. Many evaluation methods are proposed in the literature, including Agent-as-a-Judge (Zhuge et al., 2024), Mind2Web (Deng et al., 2023), etc.
However, these methods are majorly targeted for general domains and inadequate for vertical ones, especially knowledge-heavy domains like law, medicine, STEM and finance. In order to close the gap, many domain-specific benchmarks and evaluation methods are introduced, for instance LegalAgentBench (Li et al., 2024a), literature review evaluation (Tang et al., 2024), FinEval (Zhang et al., 2023a) and xbench (Chen et al., 2025). These domain-specific benchmarks and evaluations are designed to offer valuable and domain-specific insights for corresponding research.
3. FinResearchBench
3.1. The Design of FinResearchBench: an Agentic Evaluation System
Financial research is in essence a long text generation problem. And it is impractical to count on human evaluation for long text generation results due to two pitfalls, time-consuming and labor-intensive. Many LLM-as-a-Judge methods were proposed in the literature(Zheng et al., 2023; Li et al., 2024b). However, it remains questionable if directly applying LLM as a judge can robustly and reliably evaluate the results of long text generation, or more specifically in this paper, financial research agents.
The results of financial research agents are open-ended, meaning that there are no reference documents to guide the evaluation, not to mention ground truth to do a simple text-level verification. Asking experienced human experts to give a direct overall quality score is challenging enough. It seems impossible to ask a LLM, no matter how well it is prompted and aligned with humans, to give a direct score and hope it to be accurate and reliable. HelloBench(Que et al., 2024) provides category-specific checklists and asks LLMs to answer yes/no questions for each question on the given checklist. Additionally, financial research problems have distinct intricacy and subtlety, such as financial terminology, logic, cognition and deep insights. More importantly, financial research problems are widely encountered and have huge potential business impacts. Due to the strong correlation between the actual performance of financial research agents and real-world values, we are committed to providing a finance-dedicated evaluation framework featuring hierarchical structures that encompass correctness, informativeness, source attribution, professionlism, among other criteria. And the evaluation framework significantly surpasses existing methods.
3.1.1. Logic Tree
In order to draft a comprehensive and insightful professional research document in finance or any other domain, it is mandatory to do a thorough literature review, prepare ideas, and collect supporting evidence before starting to write the full report. Cognitive and linguistic theory defines different stages, i.e. Pre-Writing, Writing and Re-Writing. It defines the planning stage that contains all the preparation steps (Rohman, 1965). It lays the foundation for a potentially coherent, insightful generation outcome. In this paper, we define the underlying structure of ideas, arguments, evidences, etc. as the logic tree.
It is natural to derive that discovering the underlying logic tree of any financial research outcome and using it as the intermediate step will reduce the difficulty of the evaluation. Furthermore, the logic tree has some quantifiable metrics, such as depth, width, number of nodes, etc. Consequently, rule-based metrics can be formulated, such as analysis width, analysis depth, information density, as listed in Table 1.
3.1.2. Writing Method
It remains unclear how to extract the underlying logic tree accurately. In this section, we will discuss how professionals write in-depth research. And then it is natural to reverse the writing process and extract an accurate logic tree from a given research report.
Inspired by the cognitive writing theory (Rohman, 1965), there exists different stages like pre-writing and writing. The stage of pre-writing includes multiple operations, such as developing ideas, plans, designs; active enlistment in the cause of an idea; and consecutive logical thinking. And the stage of writing combines the ideas, evidence and words into a fresh, original, well-articulated structure. We show the process in Figure 1 and formulate it as follows:
  1. (1)
    Split or Explore: Analyze the given \$Topic\$ and decide to choose Split or Explore in a top-down manner.
    • Split: if the current \$Topic\$ is too broad or should be analyzed from different perspectives, then go with this choice. The split list of sub-topics should comply with the following requirements: (1) the sub-topics should be independent and fully cover all aspects of the source topic; (2) each sub-topic should be specific enough for further information retrieval; (3) the sub-topics should be logically related.
    • Explore: if the current \$Topic\$ is specific enough, then go with this choice to collect supporting information. A search agent will be followed up to take in the current \$Topic\$ and to retrieve relevant information. Subsequently, a well-written summary will be produced based on the \$Topic\$ and the relevant information.
  2. (2)
    Hierarchical Writing: in the order of bottom-up, perform writing for each node hierarchically. During the process, writing will be initiated on evidence nodes first, then sub-nodes, and finally the root node. The writing of a certain node will be triggered if it has no sub-node or its sub-nodes are all in ready status. This process yields a distinct argument for each topic node. Consequently, the terms topic and argument are treated as synonymous throughout this paper.
3.1.3. Logic Tree Extraction
By reversing the writing method, it is straightforward to propose the extraction method of the logic tree.
  1. (1)
    Step1: the node of the logic tree has two types: argument type and evidence type, as illustrated Figure 1.
    • Argument Type: nodes of argument type represent claims, hypotheses, or analytical insights derived from reasoning.
    • Evidence Type: nodes of evidence type represent concrete, factual statements, such as data points, statistical evidence or objectively verifiable facts.
    Leverage a well-prompted LLM to extract the research report into a logic tree, with nodes being classified into argument or evidence type and organized in a hierarchical manner. The extraction must obey the rules to preserve the original report without adding or interpreting any further information.
  2. (2)
    Step2: Leverage another well-prompted LLM to evaluate if the argument node can be supported by its subordinate argument nodes or evidence nodes.
Figure 2. The Illustration of Logic Tree Based Agent-as-a-Judge
3.1.4. Logic Tree Based Agent-as-a-Judge
Leveraging the extracted logic tree as the intermediate step, FinResearchBench is developed. It is an agentic system and essentially a logic tree based Agent-as-a-Judge, as illustrated in Figure 2. With the rich intermediate information throughout the writing process, FinResearchBench is able to provide comprehensive, reliable and robust evaluation in an automatic way. As indicated in Table 1, there are several metrics that are rule derived, either based on the extracted logic tree or not. We explain the rule-based metrics as follows. Notations used in the following section are detailed in Appendix B.
  1. (1)
    Analysis Width: The score of analysis width, $S_{width}$, is calculated based on the average number of children nodes ($\bar{N}_{child}$).
    (1) $S_{width}=\min\left(100,\max\left(0,33.33\cdot(\bar{N}_{child}-1)\right)\right)$
  2. (2)
    Analysis Depth: The score of analysis depth, $S_{depth}$, is a weighted average of the score from the tree’s maximum depth ($S_{D_{max}}$) and the score from the average depth of leaf nodes ($S_{\bar{D}_{leaf}}$).
    (2) $S_{D_{max}}$ $=\min\left(100,\max\left(0,25\cdot(D_{max}-2)\right)\right)$
    $S_{\bar{D}_{leaf}}$ $=\min\left(100,\max\left(0,40\cdot(\bar{D}_{leaf}-1.5)\right)\right)$
    $S_{depth}$ =0.$4\cdot S_{D_{max}}+6\cdot S_{\bar{D}_{leaf}}= 0$
  3. (3)
    Information Density: The score of information density, $S_{info}$, is a weighted average of the score from the total number of nodes ($S_{nodes}$) and the score from the evidence density ($S_{density}$).
    (3) $S_{nodes}$ $=\min\left(100,\max\left(0,2\cdot(N_{total}-5)\right)\right)$
[…]
    $S_{density}$ $=\min\left(100,100\cdot R_{evidence}\right)$
    $S_{info}$ =0.$7\cdot S_{nodes}+3\cdot S_{density}= 0$
  4. (4)
    Paragraph Richness: The score, $S_{rich}$, is a piecewise function of $w$, the average number of words per subtitle.
    (4) $S_{rich}(w)=\begin{cases}0&w\leq 0\\ 0.6w&0<w<100\\ 60+0.08w&200\leq w<500\\ 100&500\leq w\leq 1000\\ \max(60,100-0.05(w-1000))&w>1000\end{cases}$
  5. (5)
    Other Metrics: For dimensions that are qualitative in nature, such as Articulation Professionalism, Logic Consistency, and Viewpoint Clarity, we employ a Large Language Model (LLM) for evaluation. The LLM is prompted with a detailed rubric to assess and score each of these subjective attributes.
3.2. Data Construction
Financial research is skewed and the highly-encountered issues are closely related to monetary returns, such as financial plate investigation, stock-specific analysis, etc. Therefore, there is a huge potential for productivity improvement and real-world values for an accurate, comprehensive and reliable evaluation of these agents. Additionally, financial research agents have some unique intricacies, such as domain-specific terms, clauses and in-depth knowledge. Therefore, it is crucial to collect sufficiently complex, yet practically applicable financial research tasks.
  • Query Sampling: Randomly sample user queries from our online product: AI Cashcat (https://www.cashcat.cn). A strict privacy protection process is enforced to all raw queries before subsequent steps, including anonymization, personal identifiable information removal, session information disposal, etc.
  • Query Classification: We use a general-purpose large language model at hand (step-2 model from https://www.stepfun.com) to classify the queries into different categories. In order to determine the category list, we consult 3 domain experts, remove semantically duplicate answers and compile their responses into a union , which is: Stock-Specific Analysis, Event Analysis, Stock Selection from Given Plate, Sector Analysis, Investment Morning Brief, Post-Market Recap, Expert Opinion, Stock-Specific Performance, Stock-Specific Movements, Market Performance and Analysis, Plate Performance and Analysis, Criteria-based Stock Filtering, Trade Recommendation, Trending News, Investment Education and Others. The top 7 categories in bold font, out of the total 16 ones, occupy over 95% online traffic. Therefore, we put the focus on these 7 types of tasks.
  • Data Curation: Conduct interviews with domain experts from finance news agencies, research department of security companies, professional investors, etc., and generate 70 questions, 10 typical research problems per category.
4. Experiments and Analysis
4.1. Experiment Settings
This section outlines the experimental setup for our evaluation.
4.1.1. Core Model
We utilized Claude 3.7 Sonnet as the foundational Large Language Model (LLM) for the agentic system within the proposed FinResearchBench.
4.1.2. Agents and Experiments
We conduct a two-stage evaluation. In the first stage, we run the FinResearchBench on a range of leading international and domestic AI agents as well as our proprietary product, Cashcat Pro ( https://cashcat.cn/?mode=DeepResearch). Cashcat Pro is a deep research agent developed by FinStep, targeted specifically for the finance sector. It integrates a multi-agent framework for collaborative analysis, a suite of professional tools for real-time financial data, and cutting-edge AI models to deliver deep, actionable insights and automate the entire research process.
In the second stage, we perform a comparison between the top 3 agents versus human expert written report. We collect 20 financial research reports from reputable institutions, focusing on the types of stock-specific analysis and sector analysis. Subsequently, we extract the research topic from each of these reports and ask the top 3 agents to generate corresponding report. Finally, these research outcomes from the agents are compared with the expert written ones, under the same evaluation framework, FinResearchBench, proposed in this paper.
In this way, we obtain a holistic understanding of deep research agents in the financial domain, with not only the performance of different research agents being studied, but also the agents capability versus human experts.
4.2. Main Results
The following section presents a comprehensive evaluation of multiple deep research agents, together with expert written reports. Detailed scores are presented in Table 1 and Table 2.
Table 1. Overall Evaluation Results on FinResearchBench
Dimension Final Score
Visual-Linguistic
Synergy
Articulation
Professionalism
Articulation
Coherence
Analysis
Framework Quality
Analysis
Width
Analysis
Depth
Information
Density
Logic
Consistency
Viewpoint
Clarity
Paragraph
Richness
Weights NA 10% 10% 10% 10% 10% 10% 10% 10% 10% 10%
If Rule Derived NA No No No No Yes Yes Yes No No Yes
Gemini* 76.60 50.81 95.56 96.72 89.41 49.33 82.67 42.15 89.09 87.44 82.80
OpenAI* 74.15 24.15 88.35 89.41 84.89 53.24 79.08 48.01 90.39 93.10 90.77
Cashcat Pro 70.73 70.05 72.22 68.94 88.56 49.49 79.06 43.54 88.56 87.74 59.10
Minimax* 67.69 29.93 72.36 90.29 85.56 52.63 75.86 31.56 88.37 85.22 65.08
Reportify 67.18 36.06 70.68 50.60 86.08 56.81 79.26 51.15 87.35 85.86 67.94
Perplexity* 67.05 48.31 57.68 83.03 81.29 50.95 71.93 30.79 88.93 88.79 68.77
Genspark* 64.30 10.60 66.46 82.65 82.62 50.66 78.73 40.06 87.07 89.13 54.98
AutoGLM 59.98 0 49.06 56.82 78.31 49.35 77.05 38.16 88.56 89.13 73.26
    Not officially accessible in Chinese Mainland.
Table 2. Evaluation Results against Human Written Reports
Dimension Final Score
Visual-Linguistic
Synergy
Articulation
Professionalism
Articulation
Coherence
Analysis
Framework Quality
Analysis
Width
Analysis
Depth
Information
Density
Logic
Consistency
Viewpoint
Clarity
Paragraph
Richness
Weights NA 10% 10% 10% 10% 10% 10% 10% 10% 10% 10%
If Rule Derived NA No No No No Yes Yes Yes No No Yes
Expert Written Report 81.43 100.00 77.29 75.85 85.24 55.77 83.31 46.15 90.98 99.76 99.96
Gemini* 77.05 46.87 88.13 94.93 85.80 51.00 81.46 47.66 90.64 99.80 84.25
OpenAI* 72.03 22.13 76.40 82.47 83.07 50.29 79.36 49.82 88.58 99.87 88.27
Cashcat Pro 71.86 75.00 65.27 65.47 87.13 51.28 77.56 46.68 90.28 99.07 60.87
    Not officially accessible in Chinese Mainland.
4.2.1. Analysis of Overall Performance on FinResearchBench
Table 1 presents the overall evaluation results of various agents on FinResearchBench. The key findings are as follows:
  • Overall Ranking: Among all agents, Gemini achieves the highest Final Score (76.60), establishing as the top-performing agent in this benchmark. OpenAI is close behind with the second-highest score (74.14).
  • Gemini’s Strengths: Gemini’s leading position is predominantly driven by its outstanding performance in articulation and analysis. It scores highest in Articulation Professionalism (95.56), Articulation Coherence (96.72), Analysis Framework Quality (89.41), and Analysis Depth (82.67), indicating its strong capability to generate professional, coherent, well-structured, and in-depth financial reports.
  • OpenAI’s Strengths: OpenAI agent excels in different dimensions. It delivers the top scores in Logic Consistency (90.39), Viewpoint Clarity (93.10), and Paragraph Richness (90.77). This suggests that its outputs are highly logical, clear, and contextually rich.
  • Cashcat Pro’s Strengths: It demonstrates highly competitive performance and ranks in the top tier. It exhibits a unique, specialized skill set that differentiates it from other leading agents. Cashcat Pro shows exceptional performance in Visual-Linguistic Synergy and excels at producing highly structured and methodologically sound reports. Furthermore, among the agents accessible in Chinese Mainland, it leads by a clear margin.
4.2.2. Comparative Analysis Against Human Experts
Table 2 provides a direct comparison between the top 3 agents and expert written reports. This analysis reveals a more nuanced view of agent capabilities.
  • Human Benchmark: Expert Written Report is the leading performer with the highest Final Score (81.43), confirming that human experts still produce the highest quality.
  • Dimensions of Human Superiority: The expert’s dominance lies in dimensions requiring holistic understanding and synthesis. The expert achieves a perfect score in Visual-Linguistic Synergy (100.00) and a near-perfect score in Paragraph Richness (99.96). Additionally, the expert leads in Analysis Width (55.77), Analysis Depth (83.31), and Logic Consistency (90.98).
  • Conclusion: In summary, the evaluation results indicate that while no single AI agent surpasses human experts in terms of the overall intelligence, different agents exhibit unique strengths. The path forward to develop a superior deep research agent lies not only in improving individual metrics, but also in integrating these disparate strengths to a further extend.
4.3. Effectiveness of the Logic Tree Extraction
The accuracy of our proposed method, the logic tree based Agent-as-a-Judge, largely depends on the accuracy of the extraction of the logic tree. Thus, we design a corresponding experiment to validate the effectiveness of the logic tree extraction.
4.3.1. Objective
As the efficacy and reliability of our report evaluation methodology largely hinges upon the precision of the extracted logic tree, we aim to quantify the accuracy of the logic tree extraction using clearly defined metrics.
4.3.2. Methodology
As shown in Figure  1, the research process is composed of 2 steps: top-down Split or Explore strategy, and bottom-up Hierarchical Writing. It is intuitive to run the 2 steps in a consecutive manner and pack the outcomes of the second step and the first step to form a pair of report and ground-truth (GT) logic tree.
For each pair of report-GT logic tree, the logic tree extraction method will take over and extract corresponding logic tree from the report. And then, the similarity between the GT logic tree and the extracted one will be used to quantify the accuracy of the proposed logic tree extraction method.
4.3.3. Ground Truth Generation
We use some random queries as described in Section 3.2 to generate many pairs of report-GT logic tree. Furthermore, some rules are applied to pick out 100 pairs with the GT logic tree having different numbers of depth and width to enforce diversity on the generated dataset.
4.3.4. Similarity Metrics
We introduce 3 metrics to comprehensively compare the GT logic tree and the extracted logic tree:
  • Total Nodes Similarity: the similarity between the total number of nodes in the extracted logic tree and the GT. Here, $N_{a}$ and $N_{b}$ denote the total number of nodes in the extracted logic tree and the GT logic tree, respectively.
    $S_{\text{nodes}}=1-\frac{|N_{a}-N_{b}|}{\max(N_{a},N_{b})}$
  • Depth Similarity: the similarity between the average depths of the extracted logic tree and the GT. $d_{a}$ and $d_{b}$ are the average depths of the extracted and GT logic trees, respectively.
    $S_{\text{depth}}=1-\frac{|d_{a}-d_{b}|}{\max(d_{a},d_{b})}$
  • Width Similarity: the similarity between the average number of children (width) per node in the extracted logic tree and the GT. $w_{a}$ and $w_{b}$ refer to the average number of children per node in the extracted and GT, respectively.
    $S_{\text{width}}=1-\frac{|w_{a}-w_{b}|}{\max(w_{a},w_{b})}$
4.3.5. Results
The metrics are computed and listed as follows:
Dimension Similarity Score
Total Nodes Similarity 0.54
Depth Similarity 0.85
Width Similarity 0.76
Average Similarity 0.72
A detailed analysis indicates a significant drop in the extraction accuracy when the GT logic tree is either too large or too small. Therefore, we check the metric scores in different bins in terms of the nodes count of the GT logic tree.
Dimension [0 - 25) [25 - 50) [50, +$\infty$)∞ )
Total Nodes Similarity 0.37 0.78 0.48
Depth Similarity 0.81 0.89 0.85
Width Similarity 0.66 0.82 0.81
Average Similarity 0.61 0.83 0.71
4.3.6. Insights and Conclusions
The evaluation clearly demonstrates that the logic tree extraction is robust for moderately complex structures (25-50 nodes), with an average similarity of 0.83. However, the extraction accuracy significantly diminishes for smaller (¡25 nodes, similarity 0.61) and larger trees (¿50 nodes, similarity 0.71). This underscores the necessity to optimize the extraction step, tailored specifically towards very concise or extensively detailed reports to ensure reliability of the proposed method.
To further analyze the applicability of the logic tree, additional statistics are calculated:
Dimension [0 - 25) [25 - 50) [50, +$\infty$)∞ )
Avg. Extracted Nodes Count 37 49 48
Min. Word Count 403 10901 31946
Max. Word Count 9103 16532 50416
Spearman Correlation 0.48 (p-value: 0.0001)
We perform a correlation analysis between the total number of nodes from the extracted logic tree and the GT. As detailed in the above table, the analysis yields a Spearman correlation coefficient of $r=0.48$, with a highly statistically significant p-value of less than 0.00010.00010.0001.
This moderate positive correlation indicates that, overall speaking, an increase in the size of GT logic tree is associated with an increase in the size of the extracted logic tree. This finding supports our hypothesis that the logic tree method is sensitive to the underlying complexity of the source report.
More specifically, this relationship is particularly effective for documents up to approximately 16,000 words. As shown in the table, when the document word count increases from the first bin (max 9,103 words) to the second bin (max 16,532 words), the Avg. Extracted Nodes Count rises accordingly from 37 to 49. This demonstrates that for short to medium-length documents, our method robustly captures the increasing complexity by producing a larger and more complicated logic tree.
However, the data also suggests a potential limitation for extremely long report. As the word count extends into the third bin (over 30,000 words), the average number of extracted nodes plateaus and slightly decreases to 48. This implies that while the logic tree method serves as an accurate indicator of complexity for a majority of documents, its ability to differentiate the level of complexity diminishes for exceptionally long and elaborate reports.
5. Conclusions
5.1. Main Conclusions
In this paper, we introduce FinResearchBench, which is the first logic tree-based Agent-as-a-Judge system that is capable of performing an automatic evaluation to financial research agents and providing reliable and robust results. In addition, we collect 70 head research questions, spreading across 7 key types of tasks in the financial research domain. The problems in the domain of financial research have big potentials in investment activities, making the proposed FinResearchBench to have huge real-world values. We expect FinResearchBench to become a useful tool for the developers of any deep research agents in the financial domain and help those agents to proceed even further.
5.2. Challenges and Future Works
Despite the promising results, this study identifies several limitations that need more attention in future research:
Enhancing Multimodal Generation Capabilities: A significant gap persists in the visualization capabilities between the current AI agents and human experts. Even though current advanced agents can generate text effectively, their ability to create and integrate meaningful charts, graphs, and diagrams remain nascent. This highlights a critical direction for future work: advancing the multimodal generation capability. Future research should focus on enabling agents not only to analyze visual data but also to generate informative and aesthetically pleasing visuals that are deeply synergistic with the textual content.
Improving Logic Tree Extraction for Document of Varied Length: Our analysis indicates that the current logic tree extraction method shows varied efficacy across document of different lengths. The extraction method is suboptimal for very short and very long reports. Future work should aim to enhance the robustness of the extraction algorithm, making it less sensitive to the length of the document, and thus providing more reliable intermediate results to further improve the end-to-end evaluation.
Appendix A Question Examples
Table 3 lists some examples, one example question for one type of finance research problem.
Table 3. Question Examples
Category Example
Stock-Specific Analysis Deep research on Geely Automobile, especially the detailed assessment of the competitiveness of new vehicles and intelligent driving.
Event Analysis Please analyze the impact of the Trump administration’s proposed new tariffs on A-share listed companies within the Apple supply chain.
Stock Selection from Given Plate Conduct an in-depth analysis of the eVTOL plate, requiring comprehensive conceptual explanation and fundamental analysis of the industry to which they belong.
Sector Analysis Analyze the 2025 development trends in the cosmetics industry, identifying potential high-impact breakout opportunities and examining the commercial/industrial dynamics driving them.
Investment Morning Brief Please compile a summary of today’s investment morning brief, which should include key updates from yesterday till now, major industry events, overnight US stock market performance, and today’s opening forecasts.
Post-Market Recap Analyze the fund flows of today’s market, including the inflow and outflow of the institutional fund and the retail fund.
Expert Opinion Compile the latest opinions of the sell-side analysts over the past 12 hours and identify the three most frequently recommended stocks.
Appendix B Notations
Table 4 lists the notations used in logic tree based metrics.
Table 4. Notation for Logic Tree Based Metrics
Notation Description
$S_{width}$ The final score for Analysis Width.
$\bar{N}_{child}$ The average number of children for each node in the logic tree.
$S_{depth}$ The final score for Analysis Depth.
$S_{D_{max}}$ The score component related to the maximum depth of the tree.
$S_{\bar{D}_{leaf}}$ The score component related to the average depth of leaf nodes.
$D_{max}$ The maximum depth of the logic tree.
$\bar{D}_{leaf}$ The average depth of all leaf nodes.
$S_{info}$ The final score for Information Density.
$S_{nodes}$ The score component related to the total number of nodes.
$S_{density}$ The score component related to the density of evidence.
$N_{total}$ The total number of nodes in the logic tree.
$N_{evidence}$ The number of nodes that contain evidence.
$R_{evidence}$ The ratio of evidence-containing nodes to the total number of nodes.
$S_{rich}$ The final score for Paragraph Richness.
$w$ The average number of words per subtitle in the report.

### v3

@@
 Abstract.
-Recently, AI agents are rapidly evolving in intelligence and widely used in professional research applications, such as STEM, software development, finance, etc.
+Recently, AI agents are rapidly evolving in intelligence and are widely used in professional research applications, such as STEM, software development, and finance.
 Among these AI agents, deep research agent is a key category as it can perform long-horizon tasks and solve problems of greater complexity.
 However, there are few evaluation frameworks and benchmarks that systematically and automatically investigate the capabilities of these research agents.
-Furthermore, financial research problems have distinct complexity and subtlety.
-To fill in the gap, we propose FinResearchBench, which is a logic tree based Agent-as-a-Judge and targets specifically for the financial research agents.
+In addition, financial research problems have distinct complexity and subtlety.
+To fill in the gap, we propose FinResearchBench, which is a logic tree-based Agent-as-a-Judge and targets specifically for the financial research agents.
 It provides a comprehensive and automatic assessment of the research agents across 7 key types of tasks in the financial research domain.
-The contributions of this work are two-folded: (1) the first and innovative Agent-as-a-Judge system that extracts the logic tree of the research outcome and uses it as the intermediate information to present a comprehensive, reliable and robust evaluation; (2) finance oriented that it covers 70 typical financial research questions, spreading across 7 frequently-encountered types of tasks in the domain.
+The contributions of this work are two-folded: (1) the first and innovative Agent-as-a-Judge system that extracts the logic tree of the research outcome and uses it as the intermediate information to present a comprehensive, reliable, and robust evaluation; (2) finance-oriented that it covers 70 typical financial research questions, spreading across 7 frequently-encountered types of task in the domain.
 AI agents, financial research agents, logic tree, Agent-as-a-Judge
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generationccs: Applied computing Economics
+copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: 6th ACM International Conference on AI in Finance; November 15–18, 2025; Singapore, Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25), November 15–18, 2025, Singapore, Singaporedoi: 10.1145/3768292.3770364isbn: 979-8-4007-2220-2/2025/11
 1. Introduction
-Large Language Models (LLMs) have been rapidly developing in recent years and demonstrate superior performance in diverse tasks, such as story creation (Zhao et al., 2023), machine translation (Xu et al., 2024), dialogue system (Yi et al., 2024), sentiment analysis (Zhang et al., 2023b), etc.
+Large Language Models (LLMs) have been rapidly developing in recent years and demonstrate superior performance in diverse tasks, such as story creation (Zhao et al., 2023), machine translation (Xu et al., 2024), dialogue system (Yi et al., 2025), and sentiment analysis (Zhang et al., 2024).
 Furthermore, LLMs are dramatically evolving in both long text understanding and generation (Team et al., 2024; Anthropic, 2024).
-The intrinsic capability of LLM to generate long text is often limited, which is revealed to be bounded by the training data the model has seen during the pre-training and alignment stages (Bai et al., 2024).
-Yet, long text generation is essential, especially in professional research areas, such as literature, law, STEM and finance.
-Numerous methods have been proposed to improve LLM’s competence in long text generation, including scaling the generation length by adding long output training data (Bai et al., 2024), performing generation in a multi-agent framework (Hong et al., 2023), setting up as deep research agents (Schmidgall et al., 2025; Zheng et al., 2025), and so on.
-Deep research agent is a prominent category of AI agents and is getting more and more important, since it could generate well-articulated, source-attributed research report by autonomously browsing the internet, calling various tools, collecting information from different sources, etc.
-The research result is of high value and brings huge productivity improvement.
-Therefore, comprehensive evaluation of these deep research agents becomes critical.
-Various benchmarks and evaluation methods are proposed, such as HelloBench (Que et al., 2024), Deep Research Bench (Bosse et al., 2025), LongEval (Wu et al., 2025), etc.
+The intrinsic capability of LLM to generate long text is often limited, which is revealed to be bounded by the training data the model has seen during the pretraining and alignment stages (Bai et al., 2024).
+Yet, long text generation is essential, especially in professional research areas, such as literature, law, STEM, and finance.
+Numerous methods have been proposed to improve LLM’s competence in long text generation, including scaling the generation length by adding long output training data (Bai et al., 2024), performing generation in a multi-agent framework (Hong et al., 2024), and setting up as deep research agents (Schmidgall et al., 2025; Zheng et al., 2025).
+Deep research agent is a prominent category of AI agents and is getting more and more important, since it could generate a well-articulated, source-attributed research report by autonomously browsing the Internet, calling various tools, collecting information from different sources.
+The research result is highly valuable and brings a huge improvement in productivity.
+Therefore, a comprehensive evaluation of these deep research agents becomes critical.
+Various benchmarks and evaluation methods are proposed, such as HelloBench (Que et al., 2024), Deep Research Bench (FutureSearch et al., 2025), and LongEval (Wu et al., 2025).
 However, these methods do not fully utilize the results of the intermediate steps, which could be the deciding factor for the quality of the research outcome.
@@
 2.1. LLM and Deep Research Agent
-Large Language Models (LLMs) have shown remarkable performance in various tasks, such as story creation, machine translation, dialogue system, etc (Zhao et al., 2023; Xu et al., 2024; Yi et al., 2024; Zhang et al., 2023b).
+Large Language Models (LLMs) have shown remarkable performance in various tasks, such as story creation, machine translation, and dialogue system (Zhao et al., 2023; Xu et al., 2024; Yi et al., 2025; Zhang et al., 2024).
 In terms of dealing with long text, they also demonstrate superior capability in understanding and processing (Team et al., 2024; Anthropic, 2024).
@@
 Many AI agent systems are then developed including single-agent and multi-agent ones.
-And it represents a paradigm shift from standalone LLMs towards more autonomous, task-oriented systems (Hong et al., 2023; Wu et al., 2023).
-AI agents largely increase the upper bound of the generation results since they are able to perform more actions, like web browsing, tool calls, code execution, etc.
-Additionally, AI agents are capable of processing the obtained information more effectively and accurately, such as Chain-of-Throughts (CoT) method (Wei et al., 2022), Tree-of-Thoughts (ToT) method (Yao et al., 2023a), ReAct method (Yao et al., 2023b), etc.
-Deep research agent is a growing category of AI agents, since it can orchestrate multiple steps in an autonomous manner, including internet browsing, information retrieval, tool calling, etc.
+And it represents a paradigm shift from standalone LLMs towards more autonomous, task-oriented systems (Hong et al., 2024; Wu et al., 2023).
+AI agents largely increase the upper bound of the generation results since they are able to perform more actions, like web browsing, tool calls, and code execution.
+Additionally, AI agents are capable of processing the obtained information more effectively and accurately, such as Chain-of-Throughts (CoT) method (Wei et al., 2022), Tree-of-Thoughts (ToT) method (Yao et al., 2023a), and ReAct method (Yao et al., 2023b).
+Deep research agent is a growing category of AI agents, since it can orchestrate multiple steps in an autonomous manner, including internet browsing, information retrieval, and tool calling.
 It is able to provide in-depth research of highly complicated problems, and thus brings huge productivity improvement and high commercial values.
@@
 Moreover, the evaluation will guide the direction for further improvement.
-Numerous general benchmarks and evaluation methods are proposed, such as AgentBench (Liu et al., 2023), ToolQA (Zhuang et al., 2023), HelloBench (Que et al., 2024), DeepResearch Bench (Du et al., 2025), LongEval (Wu et al., 2025), etc.
+Numerous general benchmarks and evaluation methods are proposed, such as AgentBench (Liu et al., 2024), ToolQA (Zhuang et al., 2023), HelloBench (Que et al., 2024), DeepResearch Bench (Du et al., 2025), and LongEval (Wu et al., 2025).
 These benchmarks are constructed with diverse tools in different environments, enabling extensive evaluation of given agents.
@@
 It produces research report that contains rich information and is well articulated.
-Many evaluation methods are proposed in the literature, including Agent-as-a-Judge (Zhuge et al., 2024), Mind2Web (Deng et al., 2023), etc.
+Many evaluation methods are proposed in the literature, including Agent-as-a-Judge (Zhuge et al., 2024), and Mind2Web (Deng et al., 2023).
 However, these methods are majorly targeted for general domains and inadequate for vertical ones, especially knowledge-heavy domains like law, medicine, STEM and finance.
-In order to close the gap, many domain-specific benchmarks and evaluation methods are introduced, for instance LegalAgentBench (Li et al., 2024a), literature review evaluation (Tang et al., 2024), FinEval (Zhang et al., 2023a) and xbench (Chen et al., 2025).
+In order to close the gap, many domain-specific benchmarks and evaluation methods are introduced, for instance LegalAgentBench (Li et al., 2024a), literature review evaluation (Tang et al., 2024), FinEval (Guo et al., 2024) and xbench (Chen et al., 2025).
 These domain-specific benchmarks and evaluations are designed to offer valuable and domain-specific insights for corresponding research.
@@
 It lays the foundation for a potentially coherent, insightful generation outcome.
-In this paper, we define the underlying structure of ideas, arguments, evidences, etc.
-as the logic tree.
+In this paper, we define the underlying structure of ideas, arguments, and evidences as the logic tree.
 It is natural to derive that discovering the underlying logic tree of any financial research outcome and using it as the intermediate step will reduce the difficulty of the evaluation.
-Furthermore, the logic tree has some quantifiable metrics, such as depth, width, number of nodes, etc.
+Furthermore, the logic tree has some quantifiable metrics, such as depth, width, and number of nodes.
 Consequently, rule-based metrics can be formulated, such as analysis width, analysis depth, information density, as listed in Table 1.
@@
 Notations used in the following section are detailed in Appendix B.
+We note that the hyper-parameters in these scoring functions (e.g., constants and weights) have been determined empirically.
+This approach is taken to strike a balance between the various analytical dimensions, such as width and depth, while ensuring the resulting scores are sufficiently discriminative.
   1. (1)
@@
     $S_{\bar{D}_{leaf}}$ $=\min\left(100,\max\left(0,40\cdot(\bar{D}_{leaf}-1.5)\right)\right)$
-    $S_{depth}$ =0.$4\cdot S_{D_{max}}+6\cdot S_{\bar{D}_{leaf}}= 0$
+[…]
   3. (3)
@@
     $S_{density}$ $=\min\left(100,100\cdot R_{evidence}\right)$
-    $S_{info}$ =0.$7\cdot S_{nodes}+3\cdot S_{density}= 0$
+[…]
   4. (4)
@@
 3.2. Data Construction
-Financial research is skewed and the highly-encountered issues are closely related to monetary returns, such as financial plate investigation, stock-specific analysis, etc.
+Financial research is skewed and the highly-encountered issues are closely related to monetary returns, such as financial plate investigation and stock-specific analysis.
 Therefore, there is a huge potential for productivity improvement and real-world values for an accurate, comprehensive and reliable evaluation of these agents.
@@
   • Query Sampling: Randomly sample user queries from our online product: AI Cashcat (https://www.cashcat.cn).
-A strict privacy protection process is enforced to all raw queries before subsequent steps, including anonymization, personal identifiable information removal, session information disposal, etc.
+A strict privacy protection process is enforced to all raw queries before subsequent steps, including anonymization, personal identifiable information removal and session information disposal.
   • Query Classification: We use a general-purpose large language model at hand (step-2 model from https://www.stepfun.com) to classify the queries into different categories.
@@
 Therefore, we put the focus on these 7 types of tasks.
-  • Data Curation: Conduct interviews with domain experts from finance news agencies, research department of security companies, professional investors, etc., and generate 70 questions, 10 typical research problems per category.
+  • Data Curation: Conduct interviews with domain experts from finance news agencies, research department of security companies, professional investors, and generate 70 questions, 10 typical research problems per category.
 4. Experiments and Analysis
@@
 It integrates a multi-agent framework for collaborative analysis, a suite of professional tools for real-time financial data, and cutting-edge AI models to deliver deep, actionable insights and automate the entire research process.
+In addition, Gemini is powered by Gemini Pro 2.5, while OpenAI DeepResearch is powered by GPT-4o.
 In the second stage, we perform a comparison between the top 3 agents versus human expert written report.
@@
 AutoGLM 59.98 0 49.06 56.82 78.31 49.35 77.05 38.16 88.56 89.13 73.26
-    Not officially accessible in Chinese Mainland.
+*Not officially accessible in Chinese Mainland.
 Table 2. Evaluation Results against Human Written Reports
@@
 Cashcat Pro 71.86 75.00 65.27 65.47 87.13 51.28 77.56 46.68 90.28 99.07 60.87
-    Not officially accessible in Chinese Mainland.
+*Not officially accessible in Chinese Mainland.
 4.2.1. Analysis of Overall Performance on FinResearchBench
@@
 Therefore, we check the metric scores in different bins in terms of the nodes count of the GT logic tree.
-Dimension [0 - 25) [25 - 50) [50, +$\infty$)∞ )
+[…]
 Total Nodes Similarity 0.37 0.78 0.48
@@
 To further analyze the applicability of the logic tree, additional statistics are calculated:
-Dimension [0 - 25) [25 - 50) [50, +$\infty$)∞ )
+[…]
 Avg. Extracted Nodes Count 37 49 48
@@
 We perform a correlation analysis between the total number of nodes from the extracted logic tree and the GT.
-As detailed in the above table, the analysis yields a Spearman correlation coefficient of $r=0.48$, with a highly statistically significant p-value of less than 0.00010.00010.0001.
+As detailed in the above table, the analysis yields a Spearman correlation coefficient of $r=0.48$, with a highly statistically significant p-value of less than $0.0001$.
 This moderate positive correlation indicates that, overall speaking, an increase in the size of GT logic tree is associated with an increase in the size of the extracted logic tree.
@@
 Future work should aim to enhance the robustness of the extraction algorithm, making it less sensitive to the length of the document, and thus providing more reliable intermediate results to further improve the end-to-end evaluation.
+Moving from Logic Trees to Logic Graphs: Furthermore, a more fundamental limitation lies in the choice of a tree structure itself to represent a document’s logic.
+While trees excel at capturing hierarchical, top-down arguments, they inherently struggle to model the complex, network-like nature of sophisticated reasoning.
+Real-world argumentation is rarely a strict hierarchy; it often involves cross-references where a single piece of evidence supports multiple disparate claims, feedback loops, or premises that jointly support a conclusion in a non-linear fashion.
+A logic graph, particularly a directed acyclic graph (DAG), offers a far more robust and complete representational framework.
+Such a structure can naturally capture these many-to-many relationships, allowing for a more nuanced and accurate evaluation of a document’s overall coherence and logical integrity.
+Therefore, a pivotal direction for future work is to move beyond tree-based models and develop novel methods for extracting and analyzing these more comprehensive logic graphs.
+This would represent a significant step towards an AI that can truly understand the intricate web of arguments presented in a document.
+Acknowledgements.
+This paper was supported by ”2025 Special Program for Supporting Innovative Development in Leading Industries (AI Track) under the High-Quality Industrial Development Initiative” (Project Name: Finstep Finsmart Intelligent Service Platform; Project ID: 2025-GZL-RGZN-01024) from Shanghai Municipal Commission of Economy and Informatization, Shanghai, China.
 Appendix A Question Examples
@@
 $w$ The average number of words per subtitle in the report.
+Appendix C Example Prompt
+This appendix provides illustrative examples of the prompt used to evaluate the reports.
+C.1. Logic Tree Extraction
+The prompt used to extract the logic tree.
+Listing 1: Prompt for Logic Tree Extraction
+1"""
+2##BACKGROUND
+3Youareanexpertinlogicalanalysis,specializinginidentifyingandconstructinglogicaltreesfromtext.Toevaluatethequalityofgivenanswersbasedonmetricssuchasinformationvolume,logicalcoherence,depth,andbreadth,youmustfirstidentifythecorearguments,sub-arguments,andevidencewithinthetextandconstructalogicaltreetorepresentthem.
+4##TASK
+5Youwillreceiveaqueryandthecorrespondinggeneratedanswer.Thisanswercouldbeacasualreply,aresearchsummaryforaconcept,oraprofessionalresearchreport.Yourtaskistocarefullyreadtheanswer,identifyallcorearguments,sub-arguments,andevidence,andconstructalogicaltreetorepresenttheirrelationships.
+6Thecoreargumentservesastherootnodeofthelogicaltree,sub-argumentsareintermediatenodes,andevidenceformstheleafnodes.Allroot,intermediate,andleafnodesmustbeextracteddirectlyfromtheprovidedtext.Youmustnotaddcontentnotmentionedintheoriginaltextormakesubjectiveinferences.
+7##DEFINITIONS
+81.␣␣**Evidence**:Objectivefactspresentedinthetext,includingbutnotlimitedtostatementsofobjectivesituations,events,ordata.
+92.␣␣**Sub-argument**:Aviewpointoropinionderivedfromevidenceorothersub-argumentsthroughrigorouslogicalreasoning.
+103.␣␣**CoreArgument**:Acentralclaimderivedfromevidenceand/orsub-argumentsthatcondensestheessenceofasignificantportion,ortheentirety,ofthetext.Itrepresentsthemaintakeawayofthereport.
+11##PRINCIPLESFORCONSTRUCTINGTHELOGICALTREE
+121.␣␣**IdentifyEvidence**:Readthetextcarefullytoidentifyeverypieceofevidence.Asinglesentencemaycontainmultiplepiecesofevidence,oronepieceofevidencemaybeonlyapartofasentence.Striveforexhaustiveidentificationwithoutomission.
+132.␣␣**ConstructEvidenceNodes**:Whencreatingevidencenodes,youmust**strictlyadheretotheinformationprovidedintheoriginaltext**.Donotaddextraneousinformation,makesubjectiveinferences,oraltertheoriginalmeaning.
+143.␣␣**ConstructIntermediateNodes(Sub-arguments)**:
+15␣␣␣␣*Asub-argumentcanbeformedinoneofthreeways:
+16␣␣␣␣␣␣␣␣(i)Derivedentirelyfromevidence(leafnodes),representingaconclusiondrawndirectlyfromfacts.Thesearetypicallyfoundclosetotheleavesofthetree.
+17␣␣␣␣␣␣␣␣(ii)Derivedfromacombinationofevidenceandothersub-arguments.Thesearecommoninthemiddlelayersofthetree.
+18␣␣␣␣␣␣␣␣(iii)Derivedentirelyfromothersub-arguments,representingahigher-levelconclusion.Thesearetypicallyfoundclosertotheroot.
+19␣␣␣␣*Alldeductionsmustbelogicallysound,withaclearandexplicitrelationshipbetweenaparentnodeanditschildren.
+204.␣␣**ConstructtheRootNode(CoreArgument)**:Therootnodeisprimarilyderivedfromsub-arguments,sometimessupplementedbykeyevidence.
+21
+22UserQuery:
+23{query}
+24ArticleContent:
+25{report}
+26
+27##OUTPUTFORMAT
+28PleasestrictlyadheretothefollowingJSONformat.Note:
+291.AllstringsmustbeenclosedinEnglishdoublequotes("").
+302.Thereshouldbenoquotationmarkswithinastringvalue.
+313.EnsuretheJSONsyntaxiscorrect,includingcommas,brackets,andbraces.
+324.Donotincludeanycommentsoradditionalexplanationsintheoutput.
+335.Thetreedoesnotneedtobestrictlythreelevelsdeep;itsdepthandbreadtharedeterminedbythearticle’sstructure.
+346.Ifthearticlecontainsmultipleindependentcorearguments,therootnodecanbenull,andthecoreargumentscanbetreatedasfirst-levelchildren.
+357.Generally,thecomplexityofthelogicaltreeshouldbeproportionaltothelengthandcomplexityoftheanswer.Keepthisinmindduringextraction.
+36##REQUIREMENTS
+371.␣␣Identifythecoreargument(s)ofthetext(rootnode).
+382.␣␣Identifythesub-argumentsandevidencethatsupportthecoreargument(s)(childnodes).
+393.␣␣Statementssupportedbydataareclassifiedasevidence;allothersareconsideredarguments.
+404.␣␣Leafnodescanbeeitherargumentsorevidence,dependingonthespecificcontextandtheruleabove.
+41"""